TyroneLexWaters3rd:

Charissa is eating out with friends, and it is her turn to pay the gratuity. The bill for the food is $26.50. If she plans to tip at 20%, what should she pay if she doesn’t want to leave pennies on the table? $0.55 $1.30 $2.65 $5.30

SmokeyBrown:

Hi, and welcome to QuestionCove! To calculate the 20% tip of the $26.50 bill, we can multiply 26.50 * 0.20; the result is 5.30; this is also the same as dividing 26.50/5

TyroneLexWaters3rd:

While shopping at a clearance sale, Samantha finds a $60.00 dress on sale for 25% off. Samantha also has a 50% off coupon. Which statement correctly summarizes her savings? Samantha will save $45.00 because the two sales will combine for 75% off the original cost. Samantha will save $37.50 because she must first find the 25% sale price before taking the extra 50% reduction. Samantha will save $22.50 because she must first find the 25% sale price before taking the extra 50% reduction. Samantha will save $15.00 because the two sales will combine for 75% off the original cost.

SmokeyBrown:

In this case, I think it makes sense to first adjust the price of the dress to its 25%-off sale price and then apply the 50%-off coupon. (Actually, because both adjustments are done through multiplication, it does not matter which order we apply them; the result will be the same.) So, if the sale takes 25% off the original price, 25% of 60 is 15, so the price of the dress on sale is 60-15 = 45. (This is also the same as multiplying 60 * 0.75, since 25% off is the same as saying 75% remaining) Then, the 50% off coupon let's us divide the remaining price in half, 45/2 = 22.5. So, from the original price of 60, Samantha saves 60-22.5 = $37.50

SmokeyBrown:

This is the same result as multiplying 60*0.75*0.5 or multiplying 60*0.5*0.75; as I said, the order does not matter in this case